# Computing the Intrinsic Delaunay Triangulation of a Closed Polyhedral Surface

### v1

@@ I.1 Delaunay tessellations
Consider a closed polyhedral surface $S$, and a finite non-empty $V\subset S$ containing all singularities of $S$. We consider the definition of Bobenko and Springborn [bobenko2007discrete, Section 2] of the Delaunay tessellation of $(S,V)$. First, they define an immersed empty disk as a pair $(D,\varphi)$ where $D$ is an open metric disk of $\mathbb{R}^{2}$, and $\varphi:\overline{D}\to S$ is a map defined on the closure $\overline{D}$ of $D$ that satisfies the following: the restriction of $\varphi$ to $D$ is an isometric immersion, and $\varphi(D)\cap V=\emptyset$. Note that $\varphi$ is not necessarily injective. Bobenko and Springborn [bobenko2007discrete, Proposition 4] proved:
@@ I.2 Voronoi diagrams
Finally, consider a vertex $v$ of $\mathcal{V}$. Our third claim is that around $v$ the angles between consecutive edges of $\mathcal{V}$ are all smaller than or equal to $\pi$. To prove the claim consider the maxi-disk $(D,\varphi)$ of $v$. Let $v^{\star}$ be the center of $D$. Let $X$ be the Voronoi diagram of $\varphi^{-1}(V)$ in the plane. The faces of $X$ are all convex, being intersections of half-planes. So the angles between consecutive rays of $X$ around $v^{\star}$ are all smaller than or equal to $\pi$. There is an open disk $O$ on which $\varphi$ is injective, containing $v^{\star}$, such that $\varphi(X\cap O)=\mathcal{V}\cap\varphi(O)$. That proves the third claim, and the lemma.
@@ I.3 Voronoi diagram and Delaunay tessellation
We claim that if a directed edge $e_{0}$ of $\mathcal{V}$ is dual to a directed edge $e_{0}^{\prime}$ of $\mathcal{D}$, then the reversal of $e_{0}$ is dual to the reversal of $e_{0}^{\prime}$. The claim immediately implies the lemma, for then the duality correspondences define the desired graph isomorphism between $V$ and $\mathcal{D}$. Let us prove the claim. Let $e_{1}^{\prime}$ be the reversal of $e_{0}^{\prime}$, and let $e_{1}$ be the dual of $e_{1}^{\prime}$. We shall prove that $e_{1}$ is the reversal of $e_{0}$. We consider the maxi-disks $(D_{0},\varphi_{0})$ and $(D_{1},\varphi_{1})$ of the base vertices of $e_{0}$ and $e_{1}$, and we realize them so that they agree on the geodesic segment $p$ that is the pre-image of the common edge of $e_{0}^{\prime}$ and $e_{1}^{\prime}$. Then $\varphi_{0}$ and $\varphi_{1}$ agree on $\overline{D}_{0}\cap\overline{D}_{1}$, so they agree with a common map $\varphi_{0}\cup\varphi_{1}:\overline{D}_{0}\cup\overline{D}_{1}\to S$. Let $q$ be the geodesic segment between the centers of $D_{0}$ and $D_{1}$. Then $q$ is contained in $\overline{D}_{0}\cup\overline{D}_{1}$, and projects via $\varphi_{0}\cup\varphi_{1}$ to the common edge of $e_{0}$ and $e_{1}$ in $\mathcal{V}$. Indeed for every point $x^{\star}$ in the relative interior of $q$ the maxi-disk $(D,\varphi)$ of $\varphi(x^{\star})$ can be realized so that $x^{\star}$ is the center of $D$, and so that $\varphi$ agrees with $\varphi_{0}\cup\varphi_{1}$ on $\overline{D}\cap(\overline{D}_{0}\cup\overline{D}_{1})$. Then $\varphi^{-1}(V)$ contains exactly the two endpoints of $p$, and so $\varphi(x^{\star})$ belongs to the relative interior of an edge of $\mathcal{V}$. This proves the claim, and the lemma.
@@
Proof J.3 (Proof of Proposition 1.5, assuming Propositions J.1 and J.2).
Let $V$ contain the singularities of $\mathcal{S}(T)$, except if $\mathcal{S}(T)$ is a flat torus in which case let $V$ contain a single arbitrary vertex of $T^{1}$. Apply Proposition J.1 to replace $T$ by a portalgon $T^{\prime}$ of $O(n^{2}h)$ triangles, and to compute a subgraph $\mathcal{V}$ of $\mathcal{T}^{\prime 1}$ that is also the Voronoi diagram of $(\mathcal{S}(T),V)$, all in $O(n^{2}h\log(nh))$ time. Apply Proposition J.2 to derive from $T^{\prime}$ and $\mathcal{V}$ the portalgon of the Delaunay tessellation of $(\mathcal{S}(T),V)$ in $O(n^{2}h)$ time.
@@
Proof J.5.
Let $F$ be the face of $\mathcal{V}$ containing the faces of $T^{1}$ visited by $W$. By Lemma I.2 the face $F$ is homeomorphic to an open disk, and $w$ is the unique point of $V\cap F$. Let $\widehat{F}$ be the surface homeomorphic to a closed disk obtained by cutting the closure of $F$ along the boundary of $F$. The angles at the corners of $\widehat{F}$ are smaller than or equal to $\pi$ by Lemma I.2. So the shortest paths between those corners and $w$ are, together with the boundary edges of $\widehat{F}$, the edges of a triangulation $Y$ of $\widehat{F}$. The dual of $Y$ is a cycle, and $w$ is the central vertex of $Y$. If $\Delta$ is an unfolding of a walk in the dual of $Y$, then all occurrences of $w$ in $\Delta$ are at the same point in the plane. This easily extends to every other triangulation $Y^{\prime}$ of $\widehat{F}$, by considering a triangulation of $\widehat{F}$ that contains both $Y$ and $Y^{\prime}$.
@@
Proof J.7 (Proof of Lemma J.6).
Consider the maxi-disk $(D,\varphi)$ of $v$. Without loss of generality $\varphi$ agrees with the embedding of the neighborhood of $v$ that we fixed as a preliminary. Recall from the definition of the Delaunay tessellation that $P$ is the convex hull of $\varphi^{-1}(V)$. Let $v^{\star}$ be the middle point of $D$ (the embedding of $v$). The points in $\varphi^{-1}(V)$ correspond to the incidences between $v$ and the faces of $\mathcal{V}$ around $v$. Consider such an incidence $c$, and its corresponding point $y\in\varphi^{-1}(V)$. Consider also the face $F$ of $\mathcal{V}$ that contains $c$. The geodesic path $p$ from $v^{\star}$ to $y$ projects via $\varphi$ to a shortest path $\varphi\circ p$ from $v$ to $V$. And $\varphi\circ p$ immediately enters $F$ after leaving $v$. So the relative interior of $\varphi\circ p$ is included in $F$, and thus ends at the unique point $w\in V\cap F$. By slightly perturbing $p$ without changing its endpoints we may ensure that $\varphi\circ p$ corresponds to a walk in the dual of $T^{1}$. Then $y=\pi_{P}(x)$ by Lemma J.4. This proves the lemma.
Proof J.8 (Proof of Proposition J.2).
@@
Now, by Lemma J.6, all there remains to do is to compute, for each polygon $P$ of the portalgon, the assignment $\pi_{P}$ of positions for the vertices of $P$. Achieving the claimed linear running time when doing so requires a slight technicality. Consider a face $F$ of $\mathcal{V}$, and the point $w\in V\cap F$. Recall that for some faces $W_{0}$ of $T^{1}$ included in $F$ we need to construct a walk $W$ from $W_{0}$ to $w$ in the dual of $T^{1}$, unfold $W$, and retain the relative positions of some occurrences of $W_{0}$ and $w$ in the unfolding. Doing so independently for every face $W_{0}$ may take too long as we would visit faces of $T^{1}$ several times. Instead we consider a single spanning tree $Y$ in the dual of $T^{1}$ within $F$, we unfold the faces of $T^{1}$ that are included in $F$ along $Y$, and we retrieve all the required information from this unfolding. Note that the choice of $Y$ does not matter, and that the unfolding may overlap. Doing so for all faces $F$ of $\mathcal{V}$ takes linear time.
@@
To prove Proposition J.1 we revisit the single source shortest path algorithm described by Löffler, Ophelders, Silveira, and Staals [portalgons]. In particular we extend their algorithm to multiple sources (we let the sources be the points in $V$). The authors consider a triangulated surface, and compute the shortest paths emanating from a point $x_{0}$ on the surface by decomposing the surface according to how those paths visit the faces of the triangulation. They describe a discrete process that simulates the propagation of some waves on the surface. Their waves all start from the point $x_{0}$. In the setting of Proposition J.1, we adapt this strategy to simulate waves on $\mathcal{S}(T)$ that start from all the points in $V$, so that the waves meet along the Voronoi diagram $\mathcal{V}$ of $(\mathcal{S}(T),V)$. That simplifies the algorithm because waves now meet along a graph with geodesic edges (Lemma I.2) and do not go through singularities. We now provide a formalization of this wave algorithm. The continuous propagation of waves is discretized by a propagation of events. A crucial feature of our formalization of the algorithm is that it operates on triangles, point sets, and Voronoi diagrams in the plane $\mathbb{R}^{2}$, never in the surface $\mathcal{S}(T)$. Only the proofs of correctness will argue on the surface $\mathcal{S}(T)$. We will insist on that.
@@
Proof J.11 (Proof of Proposition J.1).
Proposition J.9 and Proposition J.10 imply that the wave algorithm can be performed in $O(n^{2}\cdot h\cdot\log(nh))$ time. Consider the returned sets $(X_{\Delta})_{\Delta}$. The sum of the cardinalities of the sets $X_{\Delta}$, summed over all the triangles $\Delta$ of $T$, is $O(n^{2}h)$ by Proposition J.9. Also, for every triangle $\Delta$, the intersection with $\Delta$ of the Voronoi diagram of $X_{\Delta}$ in $\mathbb{R}^{2}$ is the pre-image in $\Delta$ of the Voronoi diagram of $V$ in $\mathcal{S}(T)$, by Proposition J.9. Cutting each triangle $\Delta$ along the Voronoi diagram of $X_{\Delta}$, and cutting the resulting polygons into triangles along vertex-to-vertex arcs, provides the desired triangular portalgon $T^{\prime}$, along with $\mathcal{V}$.
@@
Proof J.14.
We have $V_{\Delta}\cap\overline{D}\supseteq\varphi^{-1}(V)$ by definition of the constellation $V_{\Delta}$. The other inclusion is immediate from the fact that if two immersed empty disks $(D,\varphi)$ and $(D^{\prime},\varphi^{\prime})$ are attached to $\Delta$ then $\varphi$ and $\varphi^{\prime}$ agree on $\overline{D}\cap\overline{D}^{\prime}$. Finally $\varphi^{-1}(V)\cap D=\emptyset$ by definition of an immersed empty disk (recall that $D$ is open).
Proof J.15 (Proof of Lemma J.12).
Consider a point $x\in\Delta$. There is a unique immersed empty disk $(D,\varphi)$ attached to $\Delta$ such that the center or $D$ is $x$, and such that the radius of $D$ is maximum. Then $\varphi^{-1}(V)\neq\emptyset$ because the radius of $D$ is maximum. And $\varphi^{-1}(V)=V_{\Delta}\cap\overline{D}$ by Lemma J.13. The geodesic path(s) between $x$ and the point(s) in $\varphi^{-1}(V)$ corresponds via $\varphi$ to the shortest path(s) between $\rho(x)$ and $V$. So $\rho(x)$ belongs to the Voronoi diagram of $V$ in $\mathcal{S}(T)$ if and only if $\varphi^{-1}(V)$ contains several points, equivalently $V_{\Delta}\cap\overline{D}$, which is the case if and only if $x$ belongs to the Voronoi diagram of $V_{\Delta}$ in $\mathbb{R}^{2}$.
@@
We call regions the following subsets of $\mathcal{S}(T)$: a vertex of $T^{1}$, the relative interior of an edge of $T^{1}$, and a face of $T^{1}$. The regions partition $\mathcal{S}(T)$. For every shortest path $p$ between a point $x\in\mathcal{S}(T)$ and the set $V$, record the sequence of regions intersected by $p$ when directed from $V$ to $x$. If two such paths $p$ and $p^{\prime}$ end in $\rho(\Delta)$ and have the same sequence then they correspond to the same point in the constellation $V_{\Delta}$. We claim that for every region $R$ there are $O(nh)$ sequences ending with $R$. This claim implies the lemma. Let us prove the claim. We say that a sequence is maximal if it is not a strict prefix of another sequence. And we say that a sequence is critical if it is the maximal common prefix of two distinct maximal sequences. Every critical sequence ends with a face of $T^{1}$. For every face $R^{\prime}$ of $T^{1}$ there is at most one critical sequence ending with $R^{\prime}$. Indeed every critical sequence is realized by two distinct paths. If two distinct critical sequences were to end with $R^{\prime}$, then at least two of the four associated paths would cross, and thus could be shortened, a contradiction. We proved that there are $O(n)$ critical sequences. So there are $O(n)$ maximal sequences. And every sequence contains $O(h)$ occurrences of $R$ because the happiness of $T$ is equal to $h$. This proves the claim, and the lemma.
@@
It is now convenient to name the orientation-preserving isometry of $\mathbb{R}^{2}$ that maps $s$ to $s^{\prime}$ and puts $\Delta$ side-by-side with $\Delta^{\prime}$, so let $\lambda:\mathbb{R}^{2}\to\mathbb{R}^{2}$ be this isometry. Consider the point $z=\lambda^{-1}(z^{\prime})$, and the immersed empty disk $(D,\varphi)$ attached to $\Delta$ such that the center of $D$ is $z$, and such that the radius of $D$ is maximum. Observe that $\lambda(D)=D^{\prime}$, and that $\varphi^{\prime}\circ\lambda=\varphi$. In particular the radius of $D$ is also $t$. And, crucially, $x\in\varphi^{-1}(V)$, because we already proved $x^{\prime}\in\varphi^{\prime-1}(V)$. This proves that $x\in V_{\Delta}$. And the depth of $x$ in $V_{\Delta}$ is smaller than or equal to the radius of $D$, which is $t$. The claim is proved, along with the lemma.
@@
Proof J.22.
Assuming that the geodesic segment between $x$ and $y$ intersects $\Delta$ in a point $y^{\prime}\neq y$ (at least), consider the open disk $D^{\prime}$ whose center is $y^{\prime}$ and whose boundary circle contains $x^{\prime}$. Then $D^{\prime}\subset D$. Let $\varphi^{\prime}$ be the restriction of $\varphi$ to $\overline{D}^{\prime}$. Then $(D^{\prime},\varphi^{\prime})$ is an immersed empty disk, $\varphi^{\prime}$ agrees with $\rho$ on $\Delta\cap\overline{D}^{\prime}$, and $x\in\varphi^{\prime-1}(V)$. So the depth of $x$ in $V_{\Delta}$ is smaller than or equal to the radius of $D^{\prime}$, which is smaller than the radius of $D$.
Proof J.23 (Proof of Lemma J.20).
@@
Using the claim immediately, and because $x_{0}\notin X_{\Delta_{0}}$, there is $i$ such that $x_{i}\in X_{\Delta_{i}}$ and $x_{i+1}\notin X_{\Delta_{i+1}}$, indices are modulo $k$. Consider the side $s_{i}$ of $\Delta_{i}$ that is matched to a side of $\Delta_{i+1}$. We shall prove that there is $t\leq\tau$ such that $(\tau,\Delta_{i},s_{i},x_{i})$ is a candidate event. To do so first observe that no point of $X_{\Delta_{i}}$ is closer to $y_{i}$ than $x_{i}$ because $X_{\Delta_{i}}\subseteq V_{\Delta_{i}}$ by the invariant, and because $D_{i}\cap V_{\Delta_{i}}=\emptyset$ by Lemma J.13. So $\text{Vor}(x_{i},X_{\Delta_{i}})$ intersects $s_{i}$ (at least in $y_{i}$), and its intersection with $s_{i}$ is at a distance $t\leq\tau$ from $x_{i}$ (because the distance between $y_{i}$ and $x_{i}$ is $\tau$). The tuple $(t,\Delta_{i},s_{i},x)$ is a candidate event. We are done in this case. The lemma is proved.
Proof J.24 (Proposition J.9).
@@
The wave algorithm never adds twice the same point in a set $X_{\Delta}$ of a triangle $\Delta$ of $T$. Moreover $X_{\Delta}\subseteq V_{\Delta}$ by the invariant. And the cardinality of $V_{\Delta}$ is $O(nh)$ by Lemma J.16. So the wave algorithm terminates after $O(n^{2}h)$ iterations. The algorithm does not stop until $X_{\Delta}=V_{\Delta}$ for every triangle $\Delta$ of $T$, by Lemma J.20. And the sets $(V_{\Delta})_{\Delta}$ are as desired by Lemma J.12. The lemma is proved.
@@
On the other hand if $j\in U_{i}$ is such that $\text{Vor}(z_{j},Z_{i})$ intersects $I$, then not only $\text{Vor}(z_{j},Z_{i-1})$ also intersects $I$, but $\text{Vor}(z_{j},Z_{i-1})\cap I$ contains a point in $Y$ and a point not in $Y$, so it contains the relative interior of $X_{\varepsilon}$ for some $\varepsilon\in\{0,1\}$. This proves the claim, and the lemma.

### v2

@@ I.1 Delaunay tessellations
Consider a closed polyhedral surface $S$, and a finite non-empty $V\subset S$ containing all singularities of $S$. We consider the definition of Bobenko and Springborn [4, Section 2] of the Delaunay tessellation of $(S,V)$. First, they define an immersed empty disk as a pair $(D,\varphi)$ where $D$ is an open metric disk of $\mathbb{R}^{2}$, and $\varphi:\overline{D}\to S$ is a map defined on the closure $\overline{D}$ of $D$ that satisfies the following: the restriction of $\varphi$ to $D$ is an isometric immersion, and $\varphi(D)\cap V=\emptyset$. Note that $\varphi$ is not necessarily injective. Bobenko and Springborn [4, Proposition 4] proved:
@@ I.2 Voronoi diagrams
Finally, consider a vertex $v$ of $\mathcal{V}$. Our third claim is that around $v$ the angles between consecutive edges of $\mathcal{V}$ are all smaller than or equal to $\pi$. To prove the claim consider the maxi-disk $(D,\varphi)$ of $v$. Let $v^{\star}$ be the center of $D$. Let $X$ be the Voronoi diagram of $\varphi^{-1}(V)$ in the plane. The faces of $X$ are all convex, being intersections of half-planes. So the angles between consecutive rays of $X$ around $v^{\star}$ are all smaller than or equal to $\pi$. There is an open disk $O$ on which $\varphi$ is injective, containing $v^{\star}$, such that $\varphi(X\cap O)=\mathcal{V}\cap\varphi(O)$. That proves the third claim, and the lemma. ∎
@@ I.3 Voronoi diagram and Delaunay tessellation
We claim that if a directed edge $e_{0}$ of $\mathcal{V}$ is dual to a directed edge $e_{0}^{\prime}$ of $\mathcal{D}$, then the reversal of $e_{0}$ is dual to the reversal of $e_{0}^{\prime}$. The claim immediately implies the lemma, for then the duality correspondences define the desired graph isomorphism between $V$ and $\mathcal{D}$. Let us prove the claim. Let $e_{1}^{\prime}$ be the reversal of $e_{0}^{\prime}$, and let $e_{1}$ be the dual of $e_{1}^{\prime}$. We shall prove that $e_{1}$ is the reversal of $e_{0}$. We consider the maxi-disks $(D_{0},\varphi_{0})$ and $(D_{1},\varphi_{1})$ of the base vertices of $e_{0}$ and $e_{1}$, and we realize them so that they agree on the geodesic segment $p$ that is the pre-image of the common edge of $e_{0}^{\prime}$ and $e_{1}^{\prime}$. Then $\varphi_{0}$ and $\varphi_{1}$ agree on $\overline{D}_{0}\cap\overline{D}_{1}$, so they agree with a common map $\varphi_{0}\cup\varphi_{1}:\overline{D}_{0}\cup\overline{D}_{1}\to S$. Let $q$ be the geodesic segment between the centers of $D_{0}$ and $D_{1}$. Then $q$ is contained in $\overline{D}_{0}\cup\overline{D}_{1}$, and projects via $\varphi_{0}\cup\varphi_{1}$ to the common edge of $e_{0}$ and $e_{1}$ in $\mathcal{V}$. Indeed for every point $x^{\star}$ in the relative interior of $q$ the maxi-disk $(D,\varphi)$ of $\varphi(x^{\star})$ can be realized so that $x^{\star}$ is the center of $D$, and so that $\varphi$ agrees with $\varphi_{0}\cup\varphi_{1}$ on $\overline{D}\cap(\overline{D}_{0}\cup\overline{D}_{1})$. Then $\varphi^{-1}(V)$ contains exactly the two endpoints of $p$, and so $\varphi(x^{\star})$ belongs to the relative interior of an edge of $\mathcal{V}$. This proves the claim, and the lemma. ∎
@@
Proof of Proposition 1.5, assuming Propositions J.1 and J.2.
Let $V$ contain the singularities of $\mathcal{S}(T)$, except if $\mathcal{S}(T)$ is a flat torus in which case let $V$ contain a single arbitrary vertex of $T^{1}$. Apply Proposition J.1 to replace $T$ by a portalgon $T^{\prime}$ of $O(n^{2}h)$ triangles, and to compute a subgraph $\mathcal{V}$ of $\mathcal{T}^{\prime 1}$ that is also the Voronoi diagram of $(\mathcal{S}(T),V)$, all in $O(n^{2}h\log(nh))$ time. Apply Proposition J.2 to derive from $T^{\prime}$ and $\mathcal{V}$ the portalgon of the Delaunay tessellation of $(\mathcal{S}(T),V)$ in $O(n^{2}h)$ time. ∎
@@
Proof.
Let $F$ be the face of $\mathcal{V}$ containing the faces of $T^{1}$ visited by $W$. By Lemma I.2 the face $F$ is homeomorphic to an open disk, and $w$ is the unique point of $V\cap F$. Let $\widehat{F}$ be the surface homeomorphic to a closed disk obtained by cutting the closure of $F$ along the boundary of $F$. The angles at the corners of $\widehat{F}$ are smaller than or equal to $\pi$ by Lemma I.2. So the shortest paths between those corners and $w$ are, together with the boundary edges of $\widehat{F}$, the edges of a triangulation $Y$ of $\widehat{F}$. The dual of $Y$ is a cycle, and $w$ is the central vertex of $Y$. If $\Delta$ is an unfolding of a walk in the dual of $Y$, then all occurrences of $w$ in $\Delta$ are at the same point in the plane. This easily extends to every other triangulation $Y^{\prime}$ of $\widehat{F}$, by considering a triangulation of $\widehat{F}$ that contains both $Y$ and $Y^{\prime}$. ∎
@@
Proof of Lemma J.4.
Consider the maxi-disk $(D,\varphi)$ of $v$. Without loss of generality $\varphi$ agrees with the embedding of the neighborhood of $v$ that we fixed as a preliminary. Recall from the definition of the Delaunay tessellation that $P$ is the convex hull of $\varphi^{-1}(V)$. Let $v^{\star}$ be the middle point of $D$ (the embedding of $v$). The points in $\varphi^{-1}(V)$ correspond to the incidences between $v$ and the faces of $\mathcal{V}$ around $v$. Consider such an incidence $c$, and its corresponding point $y\in\varphi^{-1}(V)$. Consider also the face $F$ of $\mathcal{V}$ that contains $c$. The geodesic path $p$ from $v^{\star}$ to $y$ projects via $\varphi$ to a shortest path $\varphi\circ p$ from $v$ to $V$. And $\varphi\circ p$ immediately enters $F$ after leaving $v$. So the relative interior of $\varphi\circ p$ is included in $F$, and thus ends at the unique point $w\in V\cap F$. By slightly perturbing $p$ without changing its endpoints we may ensure that $\varphi\circ p$ corresponds to a walk in the dual of $T^{1}$. Then $y=\pi_{P}(x)$ by Lemma J.3. This proves the lemma. ∎
Proof of Proposition J.2.
@@
Now, by Lemma J.4, all there remains to do is to compute, for each polygon $P$ of the portalgon, the assignment $\pi_{P}$ of positions for the vertices of $P$. Achieving the claimed linear running time when doing so requires a slight technicality. Consider a face $F$ of $\mathcal{V}$, and the point $w\in V\cap F$. Recall that for some faces $W_{0}$ of $T^{1}$ included in $F$ we need to construct a walk $W$ from $W_{0}$ to $w$ in the dual of $T^{1}$, unfold $W$, and retain the relative positions of some occurrences of $W_{0}$ and $w$ in the unfolding. Doing so independently for every face $W_{0}$ may take too long as we would visit faces of $T^{1}$ several times. Instead we consider a single spanning tree $Y$ in the dual of $T^{1}$ within $F$, we unfold the faces of $T^{1}$ that are included in $F$ along $Y$, and we retrieve all the required information from this unfolding. Note that the choice of $Y$ does not matter, and that the unfolding may overlap. Doing so for all faces $F$ of $\mathcal{V}$ takes linear time. ∎
@@
To prove Proposition J.1 we revisit the single source shortest path algorithm described by Löffler, Ophelders, Silveira, and Staals [22]. In particular we extend their algorithm to multiple sources (we let the sources be the points in $V$). The authors consider a triangulated surface, and compute the shortest paths emanating from a point $x_{0}$ on the surface by decomposing the surface according to how those paths visit the faces of the triangulation. They describe a discrete process that simulates the propagation of some waves on the surface. Their waves all start from the point $x_{0}$. In the setting of Proposition J.1, we adapt this strategy to simulate waves on $\mathcal{S}(T)$ that start from all the points in $V$, so that the waves meet along the Voronoi diagram $\mathcal{V}$ of $(\mathcal{S}(T),V)$. That simplifies the algorithm because waves now meet along a graph with geodesic edges (Lemma I.2) and do not go through singularities. We now provide a formalization of this wave algorithm. The continuous propagation of waves is discretized by a propagation of events. A crucial feature of our formalization of the algorithm is that it operates on triangles, point sets, and Voronoi diagrams in the plane $\mathbb{R}^{2}$, never in the surface $\mathcal{S}(T)$. Only the proofs of correctness will argue on the surface $\mathcal{S}(T)$. We will insist on that.
@@
Proof of Proposition J.1.
Proposition J.5 and Proposition J.6 imply that the wave algorithm can be performed in $O(n^{2}\cdot h\cdot\log(nh))$ time. Consider the returned sets $(X_{\Delta})_{\Delta}$. The sum of the cardinalities of the sets $X_{\Delta}$, summed over all the triangles $\Delta$ of $T$, is $O(n^{2}h)$ by Proposition J.5. Also, for every triangle $\Delta$, the intersection with $\Delta$ of the Voronoi diagram of $X_{\Delta}$ in $\mathbb{R}^{2}$ is the pre-image in $\Delta$ of the Voronoi diagram of $V$ in $\mathcal{S}(T)$, by Proposition J.5. Cutting each triangle $\Delta$ along the Voronoi diagram of $X_{\Delta}$, and cutting the resulting polygons into triangles along vertex-to-vertex arcs, provides the desired triangular portalgon $T^{\prime}$, along with $\mathcal{V}$. ∎
@@
Proof.
We have $V_{\Delta}\cap\overline{D}\supseteq\varphi^{-1}(V)$ by definition of the constellation $V_{\Delta}$. The other inclusion is immediate from the fact that if two immersed empty disks $(D,\varphi)$ and $(D^{\prime},\varphi^{\prime})$ are attached to $\Delta$ then $\varphi$ and $\varphi^{\prime}$ agree on $\overline{D}\cap\overline{D}^{\prime}$. Finally $\varphi^{-1}(V)\cap D=\emptyset$ by definition of an immersed empty disk (recall that $D$ is open). ∎
Proof of Lemma J.7.
Consider a point $x\in\Delta$. There is a unique immersed empty disk $(D,\varphi)$ attached to $\Delta$ such that the center or $D$ is $x$, and such that the radius of $D$ is maximum. Then $\varphi^{-1}(V)\neq\emptyset$ because the radius of $D$ is maximum. And $\varphi^{-1}(V)=V_{\Delta}\cap\overline{D}$ by Lemma J.8. The geodesic path(s) between $x$ and the point(s) in $\varphi^{-1}(V)$ corresponds via $\varphi$ to the shortest path(s) between $\rho(x)$ and $V$. So $\rho(x)$ belongs to the Voronoi diagram of $V$ in $\mathcal{S}(T)$ if and only if $\varphi^{-1}(V)$ contains several points, equivalently $V_{\Delta}\cap\overline{D}$, which is the case if and only if $x$ belongs to the Voronoi diagram of $V_{\Delta}$ in $\mathbb{R}^{2}$. ∎
@@
We call regions the following subsets of $\mathcal{S}(T)$: a vertex of $T^{1}$, the relative interior of an edge of $T^{1}$, and a face of $T^{1}$. The regions partition $\mathcal{S}(T)$. For every shortest path $p$ between a point $x\in\mathcal{S}(T)$ and the set $V$, record the sequence of regions intersected by $p$ when directed from $V$ to $x$. If two such paths $p$ and $p^{\prime}$ end in $\rho(\Delta)$ and have the same sequence then they correspond to the same point in the constellation $V_{\Delta}$. We claim that for every region $R$ there are $O(nh)$ sequences ending with $R$. This claim implies the lemma. Let us prove the claim. We say that a sequence is maximal if it is not a strict prefix of another sequence. And we say that a sequence is critical if it is the maximal common prefix of two distinct maximal sequences. Every critical sequence ends with a face of $T^{1}$. For every face $R^{\prime}$ of $T^{1}$ there is at most one critical sequence ending with $R^{\prime}$. Indeed every critical sequence is realized by two distinct paths. If two distinct critical sequences were to end with $R^{\prime}$, then at least two of the four associated paths would cross, and thus could be shortened, a contradiction. We proved that there are $O(n)$ critical sequences. So there are $O(n)$ maximal sequences. And every sequence contains $O(h)$ occurrences of $R$ because the happiness of $T$ is equal to $h$. This proves the claim, and the lemma. ∎
@@
It is now convenient to name the orientation-preserving isometry of $\mathbb{R}^{2}$ that maps $s$ to $s^{\prime}$ and puts $\Delta$ side-by-side with $\Delta^{\prime}$, so let $\lambda:\mathbb{R}^{2}\to\mathbb{R}^{2}$ be this isometry. Consider the point $z=\lambda^{-1}(z^{\prime})$, and the immersed empty disk $(D,\varphi)$ attached to $\Delta$ such that the center of $D$ is $z$, and such that the radius of $D$ is maximum. Observe that $\lambda(D)=D^{\prime}$, and that $\varphi^{\prime}\circ\lambda=\varphi$. In particular the radius of $D$ is also $t$. And, crucially, $x\in\varphi^{-1}(V)$, because we already proved $x^{\prime}\in\varphi^{\prime-1}(V)$. This proves that $x\in V_{\Delta}$. And the depth of $x$ in $V_{\Delta}$ is smaller than or equal to the radius of $D$, which is $t$. The claim is proved, along with the lemma. ∎
@@
Proof.
Assuming that the geodesic segment between $x$ and $y$ intersects $\Delta$ in a point $y^{\prime}\neq y$ (at least), consider the open disk $D^{\prime}$ whose center is $y^{\prime}$ and whose boundary circle contains $x^{\prime}$. Then $D^{\prime}\subset D$. Let $\varphi^{\prime}$ be the restriction of $\varphi$ to $\overline{D}^{\prime}$. Then $(D^{\prime},\varphi^{\prime})$ is an immersed empty disk, $\varphi^{\prime}$ agrees with $\rho$ on $\Delta\cap\overline{D}^{\prime}$, and $x\in\varphi^{\prime-1}(V)$. So the depth of $x$ in $V_{\Delta}$ is smaller than or equal to the radius of $D^{\prime}$, which is smaller than the radius of $D$. ∎
Proof of Lemma J.11.
@@
Using the claim immediately, and because $x_{0}\notin X_{\Delta_{0}}$, there is $i$ such that $x_{i}\in X_{\Delta_{i}}$ and $x_{i+1}\notin X_{\Delta_{i+1}}$, indices are modulo $k$. Consider the side $s_{i}$ of $\Delta_{i}$ that is matched to a side of $\Delta_{i+1}$. We shall prove that there is $t\leq\tau$ such that $(\tau,\Delta_{i},s_{i},x_{i})$ is a candidate event. To do so first observe that no point of $X_{\Delta_{i}}$ is closer to $y_{i}$ than $x_{i}$ because $X_{\Delta_{i}}\subseteq V_{\Delta_{i}}$ by the invariant, and because $D_{i}\cap V_{\Delta_{i}}=\emptyset$ by Lemma J.8. So $\text{Vor}(x_{i},X_{\Delta_{i}})$ intersects $s_{i}$ (at least in $y_{i}$), and its intersection with $s_{i}$ is at a distance $t\leq\tau$ from $x_{i}$ (because the distance between $y_{i}$ and $x_{i}$ is $\tau$). The tuple $(t,\Delta_{i},s_{i},x)$ is a candidate event. We are done in this case. The lemma is proved. ∎
Proposition J.5.
@@
The wave algorithm never adds twice the same point in a set $X_{\Delta}$ of a triangle $\Delta$ of $T$. Moreover $X_{\Delta}\subseteq V_{\Delta}$ by the invariant. And the cardinality of $V_{\Delta}$ is $O(nh)$ by Lemma J.9. So the wave algorithm terminates after $O(n^{2}h)$ iterations. The algorithm does not stop until $X_{\Delta}=V_{\Delta}$ for every triangle $\Delta$ of $T$, by Lemma J.11. And the sets $(V_{\Delta})_{\Delta}$ are as desired by Lemma J.7. The lemma is proved. ∎
@@
On the other hand if $j\in U_{i}$ is such that $\text{Vor}(z_{j},Z_{i})$ intersects $I$, then not only $\text{Vor}(z_{j},Z_{i-1})$ also intersects $I$, but $\text{Vor}(z_{j},Z_{i-1})\cap I$ contains a point in $Y$ and a point not in $Y$, so it contains the relative interior of $X_{\varepsilon}$ for some $\varepsilon\in\{0,1\}$. This proves the claim, and the lemma. ∎